課題探求法 (M) 2 単位 (選択)

Venture Business (M)

【授業目的】企業との共同研究の実施や，それを通じたベンチャービジネスへの

展開を目的とする．【授業概要】自らの研究をベンチャービジネスに発展させた場合に認定する授業

科目である．実験室内での研究に止まらず，その成果を社会に問いかける行動まで水準を高めたことは高い評価に値するという根拠による．

【授業形式】 [形態] 【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【履修要件】 [要件] 【履修上の注意】 [注意] 【到達目標】企業との共同研究や，それを通じたベンチャービジネスへ展開した

経験 · 知識を有すること．【授業計画】 [計画] 【成績評価基準】自らの研究を企業と共同研究したり，ベンチャービジネスに発展さ

せた場合，単位を認定する．打ち合わせに用いた資料等を評価の対象とする．

【教科書】 [教科書] 【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216597 【対象学生】 [対象学生] 【連絡先】 [連絡先] 【備考】 90 時間分の学修記録 (ポートフォリオ) を提出する．詳細については，指

導教員に相談，指示を仰ぐこと．

Venture Business (M) 2 units (selection)

Target⟩ Aims of the lecture are doing collaborative works with private enterprises

and establishing venture businesses through such works. Outline⟩ The lecture is aiming at founding a venture business company as well as

establishing collaborative research works with private enterprises. It’s worth to apply their accomplishment to an actual society beyond the range of a laboratory work.

Style⟩ [形態] �������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Requirement⟩ [要件] Notice⟩ [注意] Goal⟩ Objective of the lecture is to have experience and knowledge of collaborative

works with private enterprises and to mature the works to venture businesses. Schedule⟩ [計画] Evaluation Criteria⟩ Credits are ackowledged in case of doing collaborative works

with enterprises and/or maturing them to venture businesses. Handouts, reports, and presentation materials, etc. are evaluated to approve credits.

Textbook⟩ [教科書] Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216597 Student⟩ [対象学生] Contact⟩ [連絡先] Note⟩ A possible student should submit portfolios which are evidenced for his/her

90 hours’ extensive work. Students are advised to ask details about this lecture to their supervisers.